# High-Dimensional Gaussian Process Regression with Soft Kernel Interpolation

### v2

@@ B.2 Alternative Methods for Posterior Inference
Figure 5: Test RMSE of SoftKI models trained using different linear solvers on the keggundirected dataset (left) and the protein dataset (right).
Section 4.3 details the adaptation of the QR stabilized linear solve for approximate kernels (Foster et al., 2009) to the SoftKI setting. In this section, we provide empirical evidence illustrating how other linear solvers fail when confronted with datasets that generate noisy kernels. We focus on two datasets, protein and keggundirected, which have previously caused instability in Hutchinson’s pseudoloss computation.
@@
In this example we evaluate a direct solver, the Cholesky decomposition method, and CG with convergence tolerances set to $1\times 10^{-1}$, $1\times 10^{-2}$, $1\times 10^{-3}$, and $1\times 10^{-4}$. Despite adjusting the tolerances for the CG method, our experiments revealed that all solvers—except for the QR-stabilized routine—resulted in training instability. Figure 5 depicts the test RMSE performance across the different solvers. The direct and Cholesky solvers, while more stable than the CG solvers, failed to produce reliable solutions. Similarly, the CG method did not converge to acceptable solutions within the tested tolerance levels. In contrast, the QR-stabilized solver consistently produced stable and accurate solutions justifying its choice as a correctional measure that stabilizes a SoftKI on difficult problems.
@@ Appendix C Additional Experiments
We provide supplemental data for experiments in the main text (Section C.2). We also report additional comparisons with SKI-based methods (Section C.3). Finally, we provide additional data regarding SoftKI’s interpolation points (Section C.4).
@@ C.2 Supplemental to Experiments
dataset n d SoftKI SGPR SVGP
3droad 391386 3 5.586 $\pm$ 0.12 - 8.529 $\pm$ 0.216
Kin40k 36000 8 0.453 $\pm$ 0.012 0.015 $\pm$ 0.002 0.823 $\pm$ 0.01
Protein 41157 9 0.52 $\pm$ 0.006 0.015 $\pm$ 0.003 0.904 $\pm$ 0.02
Houseelectric 1844352 11 26.066 $\pm$ 1.335 - 40.33 $\pm$ 1.182
Bike 15641 17 0.227 $\pm$ 0.011 0.013 $\pm$ 0.002 0.367 $\pm$ 0.006
Elevators 14939 18 0.178 $\pm$ 0.004 0.013 $\pm$ 0.002 0.352 $\pm$ 0.046
Keggdirected 43944 20 0.554 $\pm$ 0.004 0.016 $\pm$ 0.0 0.95 $\pm$ 0.011
Pol 13500 26 0.171 $\pm$ 0.01 0.013 $\pm$ 0.002 0.325 $\pm$ 0.017
Keggundirected 57247 27 0.722 $\pm$ 0.043 0.018 $\pm$ 0.0 1.217 $\pm$ 0.047
Buzz 524925 77 7.107 $\pm$ 0.053 - 10.974 $\pm$ 0.131
Song 270000 90 3.047 $\pm$ 0.055 - 5.851 $\pm$ 0.185
Slice 48150 385 1.363 $\pm$ 0.004 0.019 $\pm$ 0.002 1.069 $\pm$ 0.014
@@
Timing. Table 6 compares the average training time per epoch in seconds to for hyperparameter optimization of SoftKI vs SVGP. We do not include SGPR since it is much faster but does not support batched stochastic optimization, and thus, is limited to smaller datasets. We observe that SVGP and SoftKI have similar performance characteristics due to the mini-batch gradient descent approach that both SVGP and SoftKI employ. SVGP requires slightly more compute since it additionally learns the parameters of a variational Gaussian distribution.
@@ C.3 Comparison with SKI-based Methods
dataset n d Exact SoftKI Skip Simplex-SKI
RMSE Protein 41157 9 0.511 ± 0.009 0.652 $\pm$ 0.012 0.817 ± 0.012 0.571 ± 0.003
Houseelectric 1844352 11 0.054 ± 0.000 0.052 $\pm$ 0.0 - 0.079 ± 0.002
@@
As we mentioned in the main text, it is not possible to apply SKI to a majority of the datasets in our experiments due to dimensionality scaling issues. Nevertheless, there are some datasets with smaller dimensionality that SKI-variants such as SKIP (Gardner et al., 2018) and Simplex-SKI (Kapoor et al., 2021) can be applied to.
@@
C.4 Additional Experiments on Interpolation Points
(a) kin40k
(b) protein
(c) bike
(d) elevators
(e) keggdirected
(f) pol
(g) keggundirected
(h) song
(i) slice
(j) 3droad
(k) buzz
(l) houseelectric
Figure 7: Interpolation and inducing points learned by each method as projected by PCA.
Figure 7 gives additional results on the PCA analysis for the UCI dataset. We observe the general trend that the interpolation points learned by SoftKI tend to be more spread out compared to the inducing points learned by SGRP and SVGP. This highlights that the structure interpolation points and inducing points are different, although they both have their start in a Nyström approximation.
@@ C.4 Additional Experiments on Interpolation Points
Figure 8: Interpolation and inducing points learned by each method as projected by UMAP.
Figure 8 visualizes the projection of the interpolation/inducing points via UMAP (McInnes et al., 2020) learned by each GP method on selected UCI datasets. As a reminder, UMAP attempts to preserve the distances between the points in the higher-dimensional manifold in the 2D plane. We make a couple of observations. First, the interpolation points learned by SoftKI are more spread out compared to the inducing points learned by SGPR and SVGP. Second, the patterns in the interpolation points learned by SoftKI confirm our intuition that high-dimensional spaces can benefit from fewer interpolation points. If a full lattice were required, then there would be less discernible structure in the projection.
@@
Figure 9: Test RMSE performance as a function of the number of interpolation points.
Figure 9 illustrates the effect of the number of interpolation points on the test RMSE of SoftKI. We find that increasing the number of interpolation points is helpful for improving test RMSE performance on some datasets (e.g., kin40k and pol) while it is detrimental for other datasets (e.g., elevators and song). In general, we expect that increasing the number of interpolation points should improve the performance of SoftKI. Thus, it is somewhat surprising that increasing the number of interpolation points did not help performance on some datasets. As a reminder, datasets such as elevators and song are challenging in general for GP models including SoftKI. Thus, we interpret the decrease in performance of SoftKI as overfitting, especially since SoftKI explains less of the data as noise compared to another interpolation point method such as SVGP.
C.5 Lengthscale and Temperature
(a) 3droad
(b) kin40k
(c) protein
(d) houseelectric
(e) bike
(f) elevators
(g) keggdirected
(h) pol
(i) keggundirected
(j) buzz
(k) song
(l) slice
Figure 10: Lengthscales.
(a) 3droad
(b) kin40k
(c) protein
(d) houseelectric
(e) bike
(f) elevators
(g) keggdirected
(h) pol
(i) keggundirected
(j) buzz
(k) song
(l) slice
Figure 11: Temperatures.

### v3

@@ Appendix B Alternative Methods for Posterior Inference
Figure 3: Test RMSE of SoftKI models trained using different linear solvers described in Section B on the keggundirected dataset (left) and the protein dataset (right).
Section 3.3 details the adaptation of the QR stabilized linear solve for approximate kernels (Foster et al., 2009) to the SoftKI setting. In this section, we provide empirical evidence that illustrates how other linear solvers fail when confronted with datasets that generate noisy kernels. We focus on two datasets, protein and keggundirected.
@@
In this example we evaluate a direct solver, the Cholesky decomposition method, and CG with convergence tolerances set to $\times 10^{-1}, 1×1$$\times 10^{-2}$, $1\times 10^{-3}$, and $1\times 10^{-4}$. Despite adjusting the tolerances for the CG method, our experiments revealed that all solvers—except for the QR-stabilized routine—resulted in training instability. Figure 3 depicts the test RMSE performance across the different solvers. The direct and Cholesky solvers, while more stable than the CG solvers, failed to produce reliable solutions. Similarly, the CG method did not converge to acceptable solutions within the tested tolerance levels. In contrast, the QR-stabilized solver consistently produced stable and accurate solutions justifying its choice as a correctional measure that stabilizes a SoftKI on difficult problems.
Dataset $n$ $d$ SoftKI m=512m{=}512 SGPR m=512m{=}512 SGPR m=512m{=}512 QR
Pol 13500 26 0.075 $\pm$ 0.002 0.127 $\pm$ 0.002 0.121 $\pm$ 0.002
Elevators 14939 18 0.36 $\pm$ 0.006 0.395 $\pm$ 0.005 0.392 $\pm$ 0.005
Bike 15641 17 0.062 $\pm$ 0.001 0.108 $\pm$ 0.004 0.105 $\pm$ 0.004
Kin40k 36000 8 0.169 $\pm$ 0.008 0.177 $\pm$ 0.004 0.169 $\pm$ 0.004
Protein 41157 9 0.596 $\pm$ 0.016 0.602 $\pm$ 0.015 0.601 $\pm$ 0.015
Keggdirected 43944 20 0.08 $\pm$ 0.005 0.099 $\pm$ 0.004 0.082 $\pm$ 0.006
Slice 48150 385 0.022 $\pm$ 0.006 0.52 $\pm$ 0.001 0.513 $\pm$ 0.001
Keggundirected 57247 27 0.115 $\pm$ 0.004 0.128 $\pm$ 0.009 0.124 $\pm$ 0.007
3droad 391386 3 0.583 $\pm$ 0.01 -- --
Song 270000 90 0.777 $\pm$ 0.004 -- --
Buzz 524925 77 0.24 $\pm$ 0.001 -- --
Houseelectric 1844352 11 0.047 $\pm$ 0.001 -- --
Table 7: Test RMSE on the UCI dataset using vanilla SGPR with SGPR using the QR stabilized linear solve described in section˜$3.$ Reference results for SoftKI are given in gray.
Since the shape of the SoftKI posterior is similar to the SGPR posterior (see Lemma A.2), we can also adapt the QR procedure used for SoftKI to SGPR. Table 7 illustrates the results of using QR solving for SGPR. We see that QR solving improves the test RMSE for SGPR. However, unlike SoftKI where QR solving is more important for training stability, we did not observe instability with an inversion of $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}+\mathbf{\Lambda}$ using the matrix inversion lemma.
@@ Appendix C Additional Experiments
We provide supplemental data for experiments in the main text (Section C.1 and Section C.2). We also report comparisons with SKI-based methods (Section C.3) and additional experimental results (Section C.4).
@@ C.1 Ricker Wavelet Experiment
Table 8: Experimental settings for the Ricker Wavlet experiment.
@@ C.2 Supplemental to Experiments
Dataset $n$ $d$ Exact GP SoftKI $m=512$ SGPR $m=512$ SVGP $m=1024$
3droad 391386 3 -- $\mathbf{5.586\pm 0.120}$ -- $8.529\pm 0.216$
[…]
Houseelectric 1844352 11 -- $\mathbf{26.066\pm 1.335}$ -- $40.330\pm 1.182$
[…]
Buzz 524925 77 -- $\mathbf{7.107\pm 0.053}$ -- $10.974\pm 0.131$
Song 270000 90 -- $\mathbf{3.047\pm 0.055}$ -- $5.851\pm 0.185$
Slice 48150 385 -- $1.363\pm 0.004$ $\mathbf{0.029\pm 0.000}$ $1.069\pm 0.014$
@@
Dataset $n$ $d$ Exact GP SoftKI m=512m{=}512 SGPR m=512m{=}512 SVGP m=1024m{=}1024
Ac-ala3-nhme 76598 126 $21.300\pm 0.502$ $1.113\pm 0.095$ $\mathbf{0.035\pm 0.000}$ $3.257\pm 0.406$
Dha 62777 168 $36.841\pm 0.375$ $1.195\pm 0.017$ $\mathbf{0.031\pm 0.000}$ $2.364\pm 0.078$
[…]
Table 10: Timing per epoch of hyperparameter optimization in seconds on MD22.
@@
Timing. Table 9 compares the average training time per epoch in seconds to for hyperparameter optimization of SoftKI vs SVGP. Note that SGPR is much faster but does not support batched stochastic optimization, and thus, is limited to smaller datasets. We observe that SVGP and SoftKI have similar performance characteristics due to the mini-batch gradient descent approach that both SVGP and SoftKI employ. SVGP requires slightly more compute since it additionally learns the parameters of a variational Gaussian distribution.
@@ C.3 Comparison with SKI-based Methods
Dataset n d Exact GP SoftKI m=512m{=}512 SKIP Simplex-SKI
RMSE
Protein 41157 9 0.511 ± 0.009 0.652 $\pm$ 0.012 0.817 ± 0.012 0.571 ± 0.003
Houseelectric 1844352 11 0.054 ± 0.000 0.052 $\pm$ 0.0 -- 0.079 ± 0.002
@@
As we mentioned in the main text, it is not possible to apply SKI to a majority of the datasets in our experiments due to dimensionality scaling issues. Nevertheless, there are some datasets with smaller dimensionality that SKI-variants such as SKIP (Gardner et al., 2018b) and Simplex-SKI (Kapoor et al., 2021) can be applied to.
@@
C.4 Lengthscale and Temperature
@@
Figure 5: Histogram of lengthscale distributions learned by automatic relevance detection (ARD) by different approximate gaussian process methods on the uci dataset.
@@
Figure 6: Histogram of Temperatures $T$ learned by SoftKI on the uci dataset.